OpenStudy (anonymous):

Hose A can fill a certain vat in 3 hours. After 2 hours of pumping, hose A is turned off. Hose B is then turned on and completes filling the vat in 3 more hours. How long would it take hose B to fill the vat alone?

OpenStudy (anonymous):

Okay this is a toughie.

OpenStudy (anonymous):

I think the answer is nine but i dont know how to set up the equation

OpenStudy (anonymous):

So hose A fills a vat in 3 hour. That means, in 2 hours, it fills 2/3 of a vat. 1/3 of a vet left and it takes hose B 3 hours to fill 1/3. 3 1/3's make a whole vat. So you multiply 3 by 3 get 9

OpenStudy (anonymous):

Sorry I don't know how to make an equation with that

OpenStudy (anonymous):

Maybe 1/9b=1

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

Solve the following for x, the time it takes b to fill the vat alone.\[\frac{1/3}{3}=\frac{1}{x}\]x = 9 hours